OpenStudy (anonymous):

\sqrt{\frac{\left(2x+2\right)}{x+2}}-\sqrt{\frac{\left(x+2\right)}{2x+2}}\ge \:\frac{7}{12}

imqwerty (imqwerty):

\(\Large\sqrt{\frac{\left(2x+2\right)}{x+2}}-\sqrt{\frac{\left(x+2\right)}{2x+2}}\ge \:\frac{7}{12}\) Lets say that \(\Large\sqrt{\frac{\left(2x+2\right)}{x+2}}=\alpha\) so we have this-> \(\Large \alpha - \frac{1}{\alpha} \ge \frac{7}{12}\) solve this and find range of \(\alpha\) \(\alpha\) is nothing but this-> \(\large \sqrt{\frac{\left(2x+2\right)}{x+2}}\) so you have the range of \(\large \sqrt{\frac{\left(2x+2\right)}{x+2}}\) you can find range of \(x\)